

********************************************************************
UARS WEEKLY STATUS REPORT - 11 JUN 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data this
past week without apparent problems.
We are evaluating the recently reprocessed V20 products before their
release to the DAAC.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to operate normally.
The HRDI team would like to say farewell and good luck to a longtime team
member, Al Marshall. Al has worked on the HRDI project since 1988 and was
responsible for writing most of the production software and keeping it
running. This was a major task and was made possible because of Al's keen
understanding of both software and scientific issues. Al was always
willing to help team members with problems with coding, operating systems,
data availability and just about everything else associated with HRDI. He
was responsible for realizing that HRDI could be operated on the dayside of
every orbit, which under the present conditions of limited spacecraft power
has significantly increased the amount of data that HRDI is collecting.
There are many other accomplishments that cannot be listed here because of
space.
Al has left us for a better place - retirement. He is one of the lucky
baby boomers (yes, a boomer!) that will be able to do so while he is still
alive. Between playing golf and fixing up his and his inlaws' houses he
should be kept busy for a while.
Al has been a major contributer to the HRDI project and his efforts are a
major reason it has been so successful. It is an unfortunate reality that
when projects like this are successful those at the top receive accolades,
while others who "toil in the trenches" are unrecognized and forgotten.
These few words are inadequate to note Al's accomplishments and
contributions, but we hope that other members of the UARS team will give
him a nod of thanks and wish him well.
***********
WINDII
***********
WINDII's next operating period is June 30 to July 9, which will
include PMC observations in the northern hemisphere. Analysis of
Version 5 processing is continuing.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
13 days CLA (7.0) PROD (L1-L3) (48R-48C=0L)
45 days WIN (5.11) PROD (L1-L3) (1540R-677-4F=859L)
The CDHF notified all PI Software Representatives concerning hardware
problems that the CDHF experienced on Disk11 last Friday. We were able to
resolve the problems prior to the weekend so production processing could
continue.
The CDHF requested the CNE re-register the CDHF5 node in the GSFC nameserver.
It is unknown how the DNS lost the node info. Any mail returned to users
stating CDHF5 is unknown, will now transfer as expected.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres.
SOLSTICE Wall Clock time increase.
"Plan to monitor Ingres file catalogue activity in debug run, to
determine where the catalogue queries are spending all of their time."
***********
FOT
***********
This report covers 29 May 1999 (Orbit 42144 , GMT Day 149) through 04 June
(GMT Day 155).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 149/0122 42144 On
WINDII 151/0124 42174 Off
PEM 152/2350 42203 AXIS On
WINDII 153/1343 42212 On
The beta angle ranged between the angles of 56.5 degrees to 27.5 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 0 degrees on 11 June 1999 (DOY 162).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.3 degrees C. Additional battery performance data
for each battery on 28 May 1999 and 04 June 1999 are provided in Appendix
D.
The clock error ranged from -10.5 to 10.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 03
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% WINDII 02
% HALOE 01
% SOLSTICE 01
% PEM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/130 - 2287.496913
Transponder B: 99/151 - 2287.499087
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/119 show the observatory data loss to be
24 hours, 13 minutes, 03.220 seconds (an increase of 0.800 seconds since
the last report period). This is a 0.0359 percent data loss which equals a
99.9641 percent data capture for the mission.
The increase of 0.800 seconds occurred over one (1) day:
7 0.800 seconds on Day 153 due to a data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The end-to-end testing has been delayed until a new FORMATS software
release can be made to accommodate the special Y2K planning aids.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL - No change
MEETINGS
FOT attended the weekly IMOC meetings on 03 June 1999.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 11
June 1999 (DOY 162).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 04 June 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
42144 149/0122 99-191 HALOE Power ON
42173 150/2347 99-193 SOLSTICE Solstice Reboot
42174 151/0124 99-192 WINDII Power OFF
42175 151/0304 99-194 OBC Clk Adj 20.6 to 0.00 msec/day
42203 152/2350 99-195 PEM AXIS Power ON
42212 153/1343 99-196 WINDII Power ON
42212 153/1414 99-197 OBC Clk Adj fr 0.00 to 20.6 msec/day
42234 155/0041 99-198 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 04 June 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
SOSFTEVLTG R-LO SOLSTICE F-TUBE VOLTAGE 42172,73
SOSGTEVLTG R-LO SOLSTICE G-TUBE VOLTAGE 42172,73
SOSNTEVLTG R-LO SOLSTICE N-TUBE VOLTAGE 42172,73
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 04 June 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
42175 151/03:10 (30 May 99) from 20.6 to 00.0 msec/day
42212 153/14:29 (02 June 99) from 00.0 to 20.6 msec/day
42234 155/00:46 (03 June 99) from 20.6 to 00.0 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 04 June 1999
28 May 1999 - GMT Day 148 Beta = 56.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -89.6/-112.0 -3.68/-4.94 +0.0/ +0.0 16.5 0.000 0.0
2 +22.4/-11.2 +5.15/+3.87 +25.6/-16.0 24.8 1.022 21.0
3 +33.6/-22.4 +4.51/+2.60 +24.8/-16.0 24.8 1.026 20.6
04 June 1999 - GMT Day 155 Beta = 27.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -89.6/-112.0 -3.05/-4.31 +0.0/ +0.0 16.5 0.000 0.0
2 +22.4/-33.6 +6.76/+5.15 +31.2/-15.2 24.0 1.022 22.5
3 +33.6/-33.6 +5.47/+4.19 +30.4/-16.4 23.8 1.022 22.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 04 June 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-028 FOT 42155 149/1823 14170 Generic Late Acquisition
99-023 FOT 42179 151/0816 14170 Generic Late Acquisition
99-028 FOT 42244 155/1702 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-027 FOT 42100 146/0249 20905 I channel dropout/no
99-023 FOT see list May 99 14170 Generic Late Acq
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-028 FOT see list June 99 14170 Generic Late Acq
SEVNTFW99/23